The ν_{7} + ν_{9} − ν_{9} Hot Band in Ethane

\underline{W. Leo Meerts}, Jos Oomens, Jörg Reuss , Department of Molecular and Laser Physics, Catholic University of Nijmegen, Toernooiveld 1, 6525ED Nijmegen, The Netherlands .

Applying an IRIR double resonance technique in a seeded molecular jet, a rotationally cooled hot band spectrum of ethane has been recorded in the region of the C–H stretching fundamental ν_{7} near 3000 cm^{−1}. A mixture of C_{2}H_{6} and SF_{6} in Ar is expanded through a slit nozzle. A 10 Watts cw CO_{2} laser excites a single rotational level in the ν_{3} manifold of SF_{6}. Collisional relaxation redistributes the excited state population over many rovibrational levels of C_{2}H_{6}. A continuously tunable F-center laser (FCL) then probes the changes in C_{2}H_{6} population induced by the CO_{2} laser. No evidence has been found for hot bands starting from excited torsional states indicating that these levels are quickly deactivated by V-T transfer to the ground state in the collisional environment of the jet. Significant population of high rotational levels in the ground state demonstrate rapid conversion of internal rotation into end-over-end rotation of the molecule. The ν_{9} state collects substantial population suggesting that this state is collisionally rather stable. The observed hot bands are assigned to the two perpendicular components of the ν_{7}+ν_{9}(A_{1g}⊕A_{2g}⊕E_{g}) ← ν_{9}(E_{u}) band. The spectrum appears to be heavily perturbed by dark state interactions in the upper level forcing us to analyse each K-stack separately. Rotational levels in the ν_{7} + ν_{9} state up to J = 15 and K = 5 have been least squares fitted to standard symmetric rotor expressions with an average standard deviation of 0.0014 cm^{−1}.